2x - 7y = -17
5x + 3y = 19
If your friend were going to solve this system of equations by first eliminating y, what general suggestions would you make so your friend could start on this in a systematic way?

Found 3 solutions by josgarithmetic, MathTherapy, ikleyn:
Answer by josgarithmetic(39618)
Answer by MathTherapy(10552)


2x - 7y = -17
5x + 3y = 19
If your friend were going to solve this system of equations by first eliminating y, what general suggestions would you make so your friend could start on this in a systematic way?
2x - 7y = - 17 ---- eq (i)
5x + 3y = 19 ------ eq (ii)

Answer by ikleyn(52788)

.

I'd recommend to this student to read the lesson
    - Solution of the linear system of two equations in two unknowns by the Elimination method
in this site.

From this lesson, the student can learn on how the method works, what is the procedure, and which particular/special
cases he (or she) can meet on the way.